OpenStudy (anonymous):

(1+2x)5=(x-2)/3+x/10

OpenStudy (anonymous):

is it \[\left( 1+2 x \right)^5=\frac{ x-2 }{ 3 }+\frac{x }{ 10 }\]

OpenStudy (anonymous):

yes but 5 will b the divisor

OpenStudy (anonymous):

\[\frac{ 1+2x }{ 5 }=\frac{ x-2}{ 3 }+\frac{ x }{ 10 }\] L.C.M. of 5,3,10=30 multiply both sides by 30 6(1+2x)=10(x-2)+3x solve for x